OpenStudy (anonymous):

little different, find the roots of the function for these two f(x) = x2 + 4 g(x) = x2 + x + 12

OpenStudy (anonymous):

FOR f(x) \[x^{2}+4=0 \rightarrow x^{2}=-4\] THUS\[x=\sqrt{-4}\]OR\[x=-\sqrt{-4}\]SO\[x=4i\]OR\[x=-4i\]

OpenStudy (anonymous):

FOR g(x) \[\Delta=1^{2}-4*(1*12)=-47\]Root1:\[\left( -1+i \sqrt{47}\right)\div 2\]Root2:\[\left( -1-i \sqrt{47}\right)\div 2\]